If 8g of a radioactive substance is present initially and 2yr later only 4g​ remain, how much of the substance will be present after 4 ​yr?
Answer #1

= 8 ek(2) Ao = 89 as t=2 A-43 A= A pkt -2k u -21 the In (0.5 - 2k In (0.5 K- 0.346573 att uyr (t=u) - (0.346513) (4) A = 8 e